A recipe calls for 3/4 cups of flour but I only want to make 2/3 how much flour should I use
Answer by jim_thompson5910(35256)

If you want to make 2/3 of the recipe (instead of the whole recipe), take 2/3 of the 3/4 cup to get (2/3)*(3/4) = (2*3)/(3*4) = 6/12 = 1/2


So instead of using 3/4 cups of flour, use 1/2 cup of flour.